### v1

# Embedding Nearly Spanning Trees

Abstract
The Erdős-Sós Conjecture states that every graph with average degree exceeding $k-1$ contains every tree with $k$ edges as a subgraph. We prove that there are $\delta>0$ and $k_{0}\in\mathbb{N}$ such that the conjecture holds for every tree $T$ with $k\geq k_{0}$ edges and every graph $G$ with $|V(G)|\leq(1+\delta)|V(T)|$.
Dedicated to the memory of Vera T. Sós.
1 Introduction
One of the best known conjectures in extremal graph theory is the Erdős-Sós Conjecture (see [Erd64]).
Conjecture 1.1 (Erdős-Sós Conjecture)
Every graph $G$ with average degree $d(G)>k-1$ contains every tree $T$ with $k$ edges as a subgraph.
Special classes of trees for which the conjecture holds include stars (this is obvious) and paths [EG59]. The conjecture also holds for large trees whose maximum degree is bounded [BPS21, Roz19]. Further, it holds if the host graph is bipartite [Ste24] or has no $4$-cycles [SW97]; for more background see [Ste21].
Instead of focussing on special classes of either trees or host graphs, a natural approach is to show the conjecture for certain ranges of $k$. Conjecture 1.1 is known to hold for $k\geq|V(G)|-4$, and Goerlich and Zak proved that it holds whenever $k\geq|V(G)|-c$, where $c$ is any given constant and $k$ is sufficiently large depending on $c$ (see [GZ16] and references therein).
We prove a much more general result, showing that there are $k_{0}$ and $\varepsilon>0$ such that for $k\geq\max\{k_{0},(1-\varepsilon)|V(G)|\}$, the Erdős–Sós conjecture holds. In other words, we show:
Theorem 1.2
There are $k_{0}\in\mathbb{N}$ and $\delta>0$ such that for all $k\geq k_{0}$ every graph $G$ with $|V(G)|\leq(1+\delta)k$ and with average degree $d(G)>k-1$ contains every tree $T$ with $k$ edges as a subgraph.
For the proof of Theorem 1.2, we need the following result by the authors.
Theorem 1.3
[RS23a, RS23b] There is an $m_{0}\in\mathbb{N}$ such that for every $m\geq m_{0}$ every graph on $m+1$ vertices that has minimum degree at least $\lfloor 2m/3\rfloor$ and a vertex of degree $m$ contains every tree T with $m$ edges as a subgraph.
Our proof of Theorem 1.2 is short and relies on the following main ideas. The high average degree of $G$ and the condition that $|V(G)|\leq(1+\delta)k$ together imply that $G$ has a subgraph $G^{\prime}$ of minimum degree at least $k-4\sqrt{\delta}k$. With this minimum degree, we can embed most of $T$ greedily into $G^{\prime}$. The problem is ensuring we can finish off the embedding. Our first step in this direction is to show that $G$ has a large set $H$ of vertices of degree at least $k$. This is done by exploiting the fact that $G$ has high average degree, but no subgraph fulfilling the conditions of Theorem 1.3.
We now distinguish two cases, according to the number of leaves of $T$. If $T$ has many leaves, we will use $H$ to embed the parents of a set $L$ of many such leaves. Because of the high degree of the vertices of $H$, we know we can leave the embedding of $L$ to the end of the process. We embed some of the more distant ancestors of $L$ into $H$, to ensure that no unembedded vertex has more than two embedded neighbours. The rest of the tree can then be embedded almost greedily into $G^{\prime}$.
If $T$ has few leaves, then it contains many vertices of degree $2$. We will embed a subset $W$ of these vertices at the very end of the process, after having already embedded their two neighbours into a pair of suitable vertices of $H$. A vertex $v\in H$ is suitable if at the time of embedding $W$, it is adjacent to most of the yet unused vertices. In other words, we need $v$ to be non-adjacent to a sufficiently large part of the used vertices. In order to achieve this for many vertices $v$, we embed most of the tree into a randomly selected set. This is possible once we have used up all vertices from $S^{\prime}:=V(G)\setminus V(G^{\prime})$. Using up $S^{\prime}$ is easily done at the beginning of the embedding process.
For all details see below.
2 The proof of Theorem 1.2
Set $\delta:=10^{-10}$. Let $m_{0}$ be given by Theorem 1.3 and set $k_{0}:=\max\{m_{0},\delta^{-2}\}.$ Let $T$ be a tree with $k\geq k_{0}$ edges, and let $G$ be a graph with $n:=|V(G)|\leq(1+\delta)k$ and $d(G)>k-1$. We can assume that $G$ is minimal with these properties. In particular, $G$ has no vertex $v$ of degree less than $\frac{k}{2}$, as deleting such a vertex would lead to a smaller graph of at least the same average degree. So,
$\delta(G)\geq\frac{k}{2},$ (1)
where $\delta(G)$, as usual, denotes the minimum degree of $G$. Set $a:=n-k$, $S\subseteq V(G)$ to be the set of all vertices of degree at most $\frac{2k}{3}+a$ (the letter $S$ stands for Small), and $b=|S|$.
If $d(v)\geq k+b$ for some $v\in V(G)$, then $G_{v}:=G\big{[}N[v]-S\big{]}$ (the subgraph of $G$ induced by $v$ and all its neighbours outside $S$) has at least $k+1$ vertices, each of which has at most $n-|V(G_{v})|\leq a-1$ neighbours outside $G_{v}$. So the minimum degree of $G_{v}$ is at least $\frac{2k}{3}+a-(a-1)>\frac{2k}{3}$. Note that $v$ has degree $|V(G_{v})|-1$ in $G_{v}$, and hence we may apply Theorem 1.3 to find that $T\subseteq G_{v}\subseteq G$, and we are done. Therefore, we assume from now on that
$\Delta(G)<k+b.$ (2)
Let $H$ be the set of all vertices of $G$ having degree at least $k$ (the letter $H$ stands for High). Then
$|H|>\frac{k}{6},$ (3)
as otherwise, (2) together with our choice of $\delta$, ensures that
$\sum_{v\in V(G)}d(v)\leq\frac{k}{6}(k+b-1)+(n-\frac{k}{6}-b)(k-1)+b(\frac{2k}{3}+a)\leq n(k-1),$
a contradiction since $d(G)>k-1$.
Observe that the number of vertices of $G$ having degree less than $k-\sqrt{ak}$ is at most $2\sqrt{ak}\leq 2\sqrt{\delta}k$ (as otherwise there are more than $ak\geq\frac{an}{2}$ non-edges and thus $d(\overline{G})>a=n-k$, a contradiction). We let $S^{\prime}$ be the set of the $\lceil 2\sqrt{\delta}k\rceil$ vertices of lowest degrees. Then for each $v\in V(G)\setminus S^{\prime}$, we have $d(v)\geq k-\sqrt{ak}$, and in particular, setting $G^{\prime}:=G[V(G)\setminus S^{\prime}]$, we have
$\delta(G^{\prime})\geq k-\sqrt{ak}-|S^{\prime}|\geq k-4\sqrt{\delta}k.$ (4)
Note that for each two vertices $u,v\in V(G^{\prime})$, we have that
$|N(u)\cap N(v)\cap V(G^{\prime})|\geq k-9\sqrt{\delta}k.$ (5)
According to whether the tree has many or few leaves, we will either use $S^{\prime}$ for embedding leaves at the end of our embedding procedure, or fill $S^{\prime}$ as early as possible. For this, we distinguish two cases.
Case 1: $T$ contains at least $10\sqrt{\delta}k$ leaves.
Choose a set $L$ of exactly $\lceil 10\sqrt{\delta}k\rceil$ leaves, and let $P_{1}$ be the set of their parents (note that not all children of $P_{1}$ need to belong to $L$). Root $T$ in an arbitrary vertex $r$ and construct a set $P_{2}$ as follows. We start by setting $P_{2}:=P_{1}\cup\{r\}$. Then, while there is a vertex $p\in V(T)\setminus P_{2}$ having at least two children in $P_{2}$, we add $p$ to $P_{2}$. If there is no such $p$, we stop the process. We note that $|P_{2}|\leq 2|P_{1}|\leq 2|L|\leq 21\sqrt{\delta}k$.
Let $P_{3}$ be the set of all parents of vertices from $P_{2}$. We embed $T[P_{2}]$ greedily into $H$, which is possible by (3) and the fact that the vertices of $H$ have degree at least $k$ in $G$ (and thus degree greater than $|P_{2}|$ into $H$). We then embed $T-P_{2}-L$ into $G^{\prime}$, going through $T-L$ in a top down fashion, starting with the root $r$, which, as it belongs to $P_{2}$, is already embedded. For each subsequent vertex $v$ that is not yet embedded, we choose its image arbitrarily among the available neighbours of the image $p_{v}$ of the parent of $v$, unless $v\in P_{3}$, in which case we embed $v$ in a vertex that is adjacent to both $p_{v}$ and the image of the unique neighbour of $v$ in $P_{2}$. This is possible by (5), because $|L|\geq 10\sqrt{\delta}k$, and since so far, we only used $G^{\prime}$ for the embedding.
It only remains to embed the vertices of $L$. Note that the parents of these vertices were embedded in $H$, and the vertices of $H$ have degree at least $k$. Thus we can embed $L$ greedily into $G$.
Case 2: $T$ has fewer than $10\sqrt{\delta}k$ leaves.
In this case, $T$ has fewer than $10\sqrt{\delta}k$ vertices of degree at least $3$. So the set $D_{2}$ of vertices of degree $2$ has size greater than $k-20\sqrt{\delta}k$. We embed the vertices of $D_{1}:=V(T)\setminus D_{2}$ arbitrarily into $G^{\prime}$ (respecting adjacencies) which is possible by (4). Let $\varphi$ denote this embedding and all future extensions of it.
Let $\mathcal{R}$ be the set of all components of $T[D_{2}]$. Note that each such component is a path and that
$|\mathcal{R}|\leq|D_{1}|\leq 20\sqrt{\delta}k.$ (6)
Take a minimal subset $\mathcal{R}^{\prime}_{1}$ of $\mathcal{R}$ such that $\bigcup\mathcal{R}^{\prime}_{1}$ contains at least $100\sqrt{\delta}k$ vertices. Choose an arbitrary path $Q$ from $\mathcal{R}^{\prime}_{1}$ and delete one of its edges, giving us two subpaths $Q_{1},Q_{2}$ of $Q$, in a way that $\mathcal{R}_{1}:=(\mathcal{R}^{\prime}_{1}\setminus\{Q\})\cup\{Q_{1}\}$ covers exactly $\lfloor 100\sqrt{\delta}k\rfloor$ vertices. Set $\mathcal{R}_{2}:=(\mathcal{R}\setminus\mathcal{R}^{\prime}_{1})\cup\{Q_{2}\}$.
For each path $R\in\mathcal{R}_{1}$, we proceed as follows. Say $R=x_{1}x_{2}\ldots x_{m}$. Set $X:=\{x_{2+3i}:0\leq i\leq m/3-1\}$ and note that $|X|\geq\lfloor 100\sqrt{\delta}k\rfloor-4|\mathcal{R}_{1}|\geq|S^{\prime}|$. We embed an arbitrary subset $X^{\prime}\subseteq X$ of size $|S^{\prime}|$ arbitrarily into $S^{\prime}$. Then we embed the vertices from $V(R)\setminus X^{\prime}$ into $G^{\prime}$, in any order. Note that at the moment of being embedded, each such vertex $v$ has at most two already embedded neighbours, at most one of which is embedded in $S^{\prime}$. Moreover, we have embedded at most $120\sqrt{\delta}k$ vertices so far. So, by (1) and (4), and as by definition $|S^{\prime}|=\lceil 2\sqrt{\delta}k\rceil$, we are able to choose an appropriate image for $v$ from a set of at least
$\frac{k}{2}-a-4\sqrt{\delta}k-120\sqrt{\delta}k>0$
vertices of $G^{\prime}$. So we can embed $\bigcup\mathcal{R}_{1}$ as planned.
Let $U\subseteq V(G)$ be the set of all vertices used so far for the embedding. Note that
$|U|\leq 120\sqrt{\delta}k\leq\frac{k}{300}.$ (7)
It remains to embed the paths from $\mathcal{R}_{2}$ into $G-U$. For this, let us introduce some notation. Given a permutation $\pi=(v_{1},v_{2},\ldots,v_{n^{\prime}})$ of $V(G)\setminus U$, we set $V_{\pi}:=\{v_{1},v_{2},\ldots,v_{\lceil\frac{49}{50}k\rceil}\}$. Let $J_{\pi}$ be the set of all indices $i<\lceil\frac{49}{50}k\rceil$ such that $v_{i}$ is not adjacent to $v_{i+1}$. Let $H_{\pi}$ be the set of all vertices in $H\setminus(U\cup V_{\pi})$ having less than $\frac{a}{3}$ non-neighbours in $G^{\prime}\setminus(U\cup V_{\pi})$.
We claim that there is a permutation $\pi$ of $V(G)\setminus U$ such that
  1. (A)
    $|J_{\pi}|\leq 30\sqrt{\delta}k$, and
  2. (B)
    $|H_{\pi}|\geq 16\sqrt{\delta}k$.
Assuming such a permutation $\pi$ exists, we can finish the embedding as we will explain now. Choose any $H^{\prime}_{\pi}\subseteq H_{\pi}$ of size exactly $\lceil 16\sqrt{\delta}k\rceil$ which is possible by (B). We start by successively embedding paths from $\mathcal{R}_{2}$ as follows until we have used all of $V_{\pi}$. We use the paths from $\mathcal{R}_{2}$ in non-decreasing order of their length. We embed each $R=x_{1}x_{2}\ldots x_{m}\in\mathcal{R}_{2}$ vertex by vertex, avoiding $H^{\prime}_{\pi}$. Say we are at vertex $x_{j}\in V(R)$ with $j\neq m$. If possible we embed $x_{j}$ in the vertex $v_{i}$ with lowest index $i$ that has not been used yet. Otherwise we can and do embed $x_{j}$ in a neighbour of $v_{i}\in V(G^{\prime})\setminus H^{\prime}_{\pi}$. Vertex $x_{m}$ has two already embedded neighbours $x,x^{\prime}$ neither of which is embedded in $S^{\prime}$, and we embed $x_{m}$ in a common neighbour of $\varphi(x)$ and $\varphi(x^{\prime})$, avoiding $H^{\prime}_{\pi}$. All of this is possible by (5), and since at any point, we have used at most $2|\mathcal{R}_{2}|+|J_{\pi}|\leq 70\sqrt{\delta}k$ vertices outside $V_{\pi}$, where the inequality holds by (6) and (A). We stop once we have used all of $V_{\pi}$, and let $R^{\prime}$ be the remainder of the path we were currently embedding. Let $\mathcal{R}_{3}$ consist of $R^{\prime}$ and all remaining paths of $\mathcal{R}_{2}$. Observe that since $|V(G)\setminus(U\cup V_{\pi})|\leq\frac{k}{50}$ and because of the order in which we used the paths from $\mathcal{R}_{2}$,
$|\mathcal{R}_{3}|\leq\frac{|\mathcal{R}_{2}|}{50}+1\leq\frac{\sqrt{\delta}k}{2}.$ (8)
where the second inequality follows from (6).
By (7), the current total amount of vertices of $G$ used for the embedding is at most $|U|+|V_{\pi}|+70\sqrt{\delta}k\leq\frac{99}{100}k-1$. Therefore,
$\big{|}\bigcup\mathcal{R}_{3}\big{|}\geq\frac{k}{100}\geq 2|H^{\prime}_{\pi}|+3|\mathcal{R}_{3}|$
and thus, there are sufficiently many vertices on the paths from $\mathcal{R}_{3}$ such that we can embed the paths from $\mathcal{R}_{3}$ as follows. For each path $x_{1}x_{2}\ldots x_{m}\in\mathcal{R}_{3}$, we successively embed all vertices $x_{j}$ with even index $j\neq m$ into $H^{\prime}_{\pi}$, as long as there still are unused vertices in $H^{\prime}_{\pi}$. For each odd index $j\notin\{1,m-1,m\}$ having the property that $x_{j-1}$ and $x_{j+1}$ are embedded in $H^{\prime}_{\pi}$, we add vertex $x_{j}$ to a set $W$, which is to be embedded at the very end. Observe that by construction, and by (8),
$|W|\geq|H^{\prime}_{\pi}|-|\mathcal{R}_{3}|\geq 15\sqrt{\delta}k.$
We now use (5) to embed all remaining vertices from $V(\bigcup\mathcal{R}_{3})\setminus W$ into $G^{\prime}$.
Finally, we embed $W$. By construction, each vertex of $W$ is an $x_{j}$ from some path of $\mathcal{R}_{3}$, with $x_{j-1},x_{j+1}$ embedded in vertices $u,v\in H^{\prime}_{\pi}\subseteq H_{\pi}$. By definition of $H_{\pi}$, at most $\frac{2a}{3}$ vertices in $V(G)\setminus(U\cup V_{\pi})$ are not common neighbours of $u$ and $v$. So, as $|V(T)|=n-a+1$ and $U\cup V_{\pi}$ has been used, we are able to find a common neighbour of $u$ and $v$ in which to embed $x_{j}$. This finishes the embedding of $T$.
It only remains to prove our claim that there is a permutation of $V(G)\setminus U$ such that (A) and (B) hold. We take a random permutation $\pi=(v_{1},v_{2},\ldots,v_{n^{\prime}})$ of $V(G)\setminus U$, and show that with positive probability, it has both these properties. We note that by (4), and since $V(G)\setminus U\subseteq V(G^{\prime})$, we have
$\mathbb{E}\big{[}|J_{\pi}|\big{]}\leq\sum_{v\in V_{\pi}}\frac{|V(G)\setminus N(v)|}{|V_{\pi}|}\leq|V_{\pi}|\cdot\frac{4\sqrt{\delta}k+a}{|V_{\pi}|}\leq 5\sqrt{\delta}k.$
Hence by Markov’s inequality (see [MR02]), the probability that (A) fails is at most $\frac{1}{6}$.
We will show that (B) fails with probability less than $\frac{5}{6}$, which will finish the proof of our claim. By definition of $H$, each vertex from $H$ has less than $a$ non-neighbours in $V(G^{\prime})\setminus U$. Moreover, as $|U|\geq|S^{\prime}|>a$, for any $v\in V(G^{\prime})\setminus U$, the probability that $v\notin V_{\pi}$ is at most $\frac{1}{50}$. So, setting
$s_{\pi}:=\sum_{u\in H\setminus(U\cup V_{\pi})}|\{v\ |\ u\notin U,uv\notin E(G^{\prime})\}\setminus V_{\pi}|,$
we have that
$\mathbb{E}[s_{\pi}]\leq\frac{|H\setminus U|a}{2500}\leq\frac{ak}{2400},$
and by Markov’s inequality (see [MR02]), the probability that $s_{\pi}\geq\frac{ak}{1600}$ is at most $\frac{2}{3}$. In particular, the probability that $H\setminus(U\cup V_{\pi})$ has more than $\frac{k}{500}$ vertices which each have at least $\frac{a}{3}$ non-neighbours in $V(G^{\prime})\setminus(U\cup V_{\pi})$ is at most $\frac{2}{3}$. So, if we can show that the probability that $|H\setminus(U\cup V_{\pi})|<\frac{k}{500}+16\sqrt{\delta}k$ is at most $\frac{1}{6}$, we are done.
For this, note that $|H\setminus U|>\frac{4}{25}k$ by (3) and (7). Also by (7), for each $v\in V(G^{\prime})\setminus U$, the probability that $v\notin V_{\pi}$ is at least $\frac{1}{60}$. It follows that the expectation of $|H\setminus(U\cup V_{\pi})|$ is at least $\frac{k}{375}$. Applying the Chernoff bound (see [MR02]), we deduce that the probability that $|H\setminus(U\cup V_{\pi})|<\frac{k}{400}$ is (much) less than $\frac{1}{6}$. As $\frac{k}{400}>\frac{k}{500}+16\sqrt{\delta}k$, we are done.